OpenStudy (anonymous):

The square root of absolute value of 16 divided by the absolute value of 100?

OpenStudy (anonymous):

Is this what your question looks like? \(\Large{\sqrt{\frac{|16|}{|100|}}}\)

OpenStudy (anonymous):

Yes kinda yea

OpenStudy (lncognlto):

= sqr (16/100) = + or - (4/10) = + or - (2/5)

OpenStudy (anonymous):

Assuming I'm understanding the question correctly, your answer of \(\pm\frac{2}{5}\) is correct.